### v1

# Precomputed Dominant Resource Fairness

Abstract
Although resource allocation is a well studied problem in computer science, until the prevalence of distributed systems, such as computing clouds and data centres, the question had been addressed predominantly for single resource type scenarios. At the beginning of the last decade, with the introuction of Dominant Resource Fairness, the studies of the resource allocation problem has finally extended to the multiple resource type scenarios. Dominant Resource Fairness is a solution, addressing the problem of fair allocation of multiple resource types, among users with heterogeneous demands. Based on Max-min Fairness, which is a well established algorithm in the literature for allocating resources in the single resource type scenarios, Dominant Resource Fairness generalises the scheme to the multiple resource case. It has a number of desirable properties that makes it preferable over alternatives, such as Sharing Incentive, Envy-Freeness, Pareto Efficiency, and Strategy Proofness, and as such, it is widely adopted in distributed systems. In the present study, we revisit the original study, and analyse the structure of the algorithm in closer view, to come up with an alternative algorithm, which approximates the Dominant Resource Fairness allocation in fewer steps. We name the new algorithm Precomputed Dominant Resource Fairness, after its main working principle.
Keywords Resource Distribution  $\cdot$ Dominant Resource Fairness  $\cdot$ Precomputed
1 Introduction
Dominant Resource Fairness (DRF) is a resource allocation scheme, which aims to provide fairness of allocation among users with heterogeneous demands, in a time-sharing system with multiple resource types, such as computing clouds (Ghodsi et al., 2011). The services offered by these systems may include processing power, memory, and bandwidth, all of which are expected to be distributed fairly among the clients, according to some definition of fairness.
Although the problem has been well-studied for single resource types (Keshav, 1997; Bertsekas et al., 1992), prior to DRF, less attention was paid to the multi-resource scenarios. The general approach was to adopt well established solutions for single resource allocation, such as Max-min Fairness (MF) and its derivatives, for a fixed bundle with differing amounts of resources from each resource type; e.g. standardised virtual machines. But this approach leads to inefficiencies, since the needs of different users are heterogeneous, where one needs predominantly one resource type, the other needs another. In the end, users ended up with unused resources that come in the bundle with the resources they need, where at the same time there were other users in the system in demand of those unused resources.
Built on the MF premise, which aims to maximise the minimum share given to any user in a single resource setting, DRF aims to generalise the scheme by maximising the minimum relative share given to any user, over her dominant share; which in turn is defined for each user as the highest ratio of her demanded resources to their total reserve in the system (e.g. the ratio of demanded memory to the total available memory). The rest of the resources is allocated, for each user, in their fixed proportions to the user’s dominant share, which is known as the Leontief Preferences in the economy literature.
DRF has acquired wide adoption in the distributed systems and cluster computation because of its properties, such as being Pareto optimal, Envy Free, Strategy Proof, and creating Sharing Incentive. In the present paper, we analyse the DRF algorithm as introduced by Ghodsi et al. (2011) in closer view, and based on the observations gathered thereby, introduce a new algorithm to approximate the DRF allocation in fewer steps. We call this new algorithm Precomputed Dominant Resource Fairness (PDRF), after its main working principle of precomputing the main allocation loop of DRF, and assigning resources to users without going through the tedious iterations of the loop.
One study that also needs to be mentioned at the onset, is that of Parkes et al. (2015), the method and the findings of which closely resemble to that of the present study, with some key differences that allow us to implement our newly proposed algorithm. The analysis method and the algorithm they propose, the Extended Dominant Resource Fairness (EDRF) are similarly based on precomputing the final allocation of DRF, but they modeled the underlying linear programming problem, rather than the algorithm itself. They construct a theoretical framework that allows them to provide more rigorous proofs of DRF properties, which is the main objective of the article. We elaborate on this in Section 3.2.
The rest of this paper is organised as follows: In the next section, we briefly review the literature for related work. In Section 3.4, we present our model, after reviewing DRF and EDRF in subsections 3.1 and 3.2, respectively, and introducing an analysis method in subsection 3.3. In section 4, we present the results of the tests we carried out with PDRF. In Section 5 we analyse PDRF performance. Section 6 consists of three subsections, in which we discuss higher order cycles in PDRF (6.1), Pareto Efficiency property (6.2), and weighting policies (6.3), which we believe, are important points to focus on critically. Finally, we conclude in Section 7.
2 Related Work
DRF is first proposed by Ghodsi et al. (2011), where they introduce the new notion of establishing an MF allocation among dominant shares, rather than a fixed resource bundle of different resource types. They also provide an algorithm to carry out the distribution, and produce empirical data to demonstrate the performance of the algorithm. In their theoretical sections, they compare the virtues of the newly introduced allocation scheme to the ones existing in the economy literature, for this is as much a problem in the field of economy as it is in the field of computer science. Their results clearly demonstrate that DRF possesses various advantages over its alternatives.
The model is challenged by Dolev et al. (2012). They developed an alternative fairness notion, which they refer to as no justified complaints, based on the assumption that the fairness of distribution of resources should be assessed from a global system point of view, considering the bottleneck property of resources. A resource is a bottleneck resource, if the availability of this resource is limited and there is a race condition for accessing it. They developed the Bottleneck Based Fairness (BBF) algorithm to implement this paradigm as an alternative to DRF, with relative advantages of each algorithm in different use cases.
Not only challenged, but DRF has also been adopted, generalised, contextualised, or extended to different cases. For example, Gutman and Nisan (2012), introduced an economical framework to compare and contrast the two systems described above. While doing so they also critically assessed both approaches from an economic point of view. Joe-Wong et al. (2013) investigates the trade-off between fairness and efficiency, providing a unifying framework with two families of fairness functions, DRF being a special case in one of them. Kash et al. (2014) takes a scenario in which the complete information on demands is not available, since users arrive and leave the system, and develops a Dynamic DRF (DDRF) to address the problem. The DDRF algorithm is improved for its performance in Li et al. (2015).
Finally, DRF has been reformulated by Parkes et al. (2015) to EDRF, in order to study its properties more rigorously. EDRF is also a solution to a special case, where the resources, as well as user demands are well-divisible, i.e. fractions of demands can be allocated to users, but for obvious reasons this is not widely applicable to real life scenarios, where half a task does not provide any added utility to the users. In the same article, they also introduce Sequential Min-max Algorithm to address the indivisible demands scenario. One more contribution of EDRF is that it addresses the situations where there are users who submit zero demands for one or more resources.
We now proceed with presenting DRF and EDRF models, as well as and in relation to our own, in the following section.
3 The Model
We start this section by briefly presenting DRF, and EDRF, and then proceed to introducing PDRF. The notions we introduce in the preceding subsections will help us to construct our own model.
3.1 Dominant Resource Fairness
The DRF allocation scheme is build on the concept of dominant shares. The dominant share for each user is defined as the maximum of the fraction of her demands for different resource types to their respective reserves in the system. As exemplified in Ghodsi et al. (2011), in a system with $9$ CPUs and $18$ GB memory, the dominant share, $ds$ for user A, with the demand vector $\left\langle 1,4\right\rangle$ is memory, since the fractional demands, $fd$, of the user are $1/9$ and $4/18=2/9$, respectively. Similarly, the $ds$ of a user with the demand vector $\left\langle 3,1\right\rangle$ is for CPU, since $fd$’s are $3/9$ and $1/18$, respectively. The question then is, what is a fair allocation between the users, whose demands for the resources are defined as such.
If the number of tasks allocated to users A and B is denoted by $x$ and $y$, respectively, the DRF allocation is defined as the solution to the optimising problem below (Ghodsi et al., 2011, p. 4):
$\max(x,y)$ Maximise allocation
$x+3y$$\leq$ 99 99 CPU Constraint
$4x+y$$\leq$ 1818 1818 Memory Constraint
$\frac{2x}{9}$ $=$ $\frac{y}{3}$ Equalise Dominant Shares
Designed to solve this problem, DRF is a progressive filling algorithm, applying MF on the users’ $ds$’s by looping on the total allocated $ds$’s of users, taking the least allocated at each iteration, and allocating one unit of her demand vector, assumed to correspond to one task, until one of the resources is exhausted. Although not explicitly stated, in addition to this mechanism, DRF utilises a tie-breaking policy to take the user with the higher $ds$, in instances when more than one user is allocated the same minimum share. This we see at the example given in the article, in the allocation process for the above mentioned case.
In the initial state, where both user’s allocated $ds$ is 00, the algorithm starts assigning one task to user B, which updates the state to 00 allocated $ds$ for A, and $1/3$ for B. Since now the least allocated $ds$ belongs to A, the algorithm picks her and assigns one task, updating the state to $2/9$ w.r.t. $1/3=3/9$. Still A has the lower allocated $ds$, so one more task is allocated to her, updating to $4/9$ w.r.t. $3/9$, and pushing B to the position of the user with the least allocated $ds$. Thusly the algorithm proceeds with B, and continuing further in the similar fashion, assigns $2/3$ dominant shares to each user, at which point the CPU is saturated and the algorithm stops.
As seen in the example above, DRF is a progressive filling algorithm, since it aims to keep the growth of allocation among different agents at the same pace. DRF reduces to MF, and also to BBF (Dolev et al., 2012), in cases where all users’ dominant resources are the same.
3.2 Extended Dominant Resource Fairness
The study by Parkes et al. (2015) introduces an alternative view to DRF. Instead of an iterative algorithm, they aim to precompute the resulting distribution, with a linear program. For doing so, they employ two notions in the setting. First, in a given user demand vector, the ratio between the entries for different resources is fixed. Second, the operating principles of progressive filling algorithm dictates that the pace at which the allocation for dominant shares of different users grow should be equated. Departing from these notions, it is possible to normalise each demand vector to its dominant share, and abstract that they will all be allocated simultaneously. In other words, if we scale all the dominant shares to $1$, scale all other fractional shares accordingly, we end up in a setting in which at each step, every user takes $1$ normalised allocation, and fairness among the dominant shares is kept intact.
The question then is, what is the maximum number of such allocations that can be made before depleting one of the resources? The answer starts with normalising the vector of fractional demands,$fd_{ir}$ to the dominant share, $ds_{i}$, which is denoted $d_{ir}$ and defined more explicitly as:
$d_{ir}=\frac{fd_{ir}}{ds_{i}}$
As such, each user’s dominant share is equated to $1$ and the question is reduced to find the maximum scaling factor $x$ such that:
$\sum_{i=1}^{n}x\cdot d_{ir}\leq 1,\forall r\in R$
where $R$ is the set of resources. The answer is obtained simply by extracting $x$ out, and passing the summation to the other side of the inequality; obtaining:
$x=\frac{1}{\max_{r}\sum_{i}d_{ir}}$
At this point, the authors allow themselves a rather unrealistic abstraction of both well-divisible resources and well-divisible tasks. But although unrealistic, it is not unreasonable, since the abstraction lends a great ease for the mathematical analysis of DRF properties, without loss of generality.
Before explicitly describing the exact allocation of each resource to each user, they go on to introduce a weighting mechanism, and then suffice to describe the assignment process in the pseudocode for the weighted version, of which the unweighted version is a special case where all the users are weighted equally for all the resources. Yet, it is rather straightforward that the allocation of resource $r$ to user $i$ is given by $x\cdot d_{ir}\cdot r$.
The rest of the article is rather notationwise involved and out of the scope of the present study, since we are not investigating the DRF properties, but taking them as they are in the literature, and aiming to improve on the algorithm. They also present another algorithm for dropping the well-divisibility assumption for tasks, but it is an iterative algorithm and precomputing is not involved, leaving us no bases for comparison. Thus, in the present study, we will suffice to present the model to this extent.
One important contribution of the article is to point out the fact that the resulting distribution cannot guarantee Pareto efficiency, as such. In order for an allocation to be Pareto efficient, it should not be possible to increase the utility of one user without decreasing another. However, the algorithm stops when a resource is depleted; yet, there may be tasks who do not need the depleted resource, and the remaining resources may be sufficient to further increase those tasks’ utilisation, without decreasing any other task’s. To address this issue, the algorithm takes multiple allocation rounds, each one ending with the depletion of one resource, and the next one starting after removing the saturated demands, until all demands are saturated.
3.3 Cyclic Structure of the DRF Main Loop
We will shortly move on to introducing PDRF, but before doing so, we will present a key observation that led to the development of the algorithm. It also introduces, what we believe is a simpler and more intuitive approach to precomputing the resulting DRF allocation for given sets of demands and resources.
As indicated, the progressive filling algorithm, which DRF is based upon, operates on the principle of keeping the rate of growth among the tasks of each user at the same pace. As a consequence, the number of iterations at which a given task is allocated one unit of its demand vector is inversely proportional to its dominant share, the tasks with lower dominant shares being visited more often, where the task with the highest dominant share, $ds^{*}$, is visited the least.
Departing from this observation, it is possible to identify cycles within the main allocation loop. The structure of the cycle is dependent on the distribution of the demands and the preferred tie-breaking policy in case of equal dominant shares. For example, the original DRF algorithm takes the demand with $ds^{*}$ first, in cases where the users have been allocated the same amount of resources thereto, as seen in the example in Section 3.1. The user with $ds^{*}$, User B, with $ds=1/3$, takes the first iteration, and two iterations of the User A with $ds=2/9$ completes the cycle. One more iteration of the cycle exhausts the reserve of the available CPU, and thus completes the main loop.
The exact length and the ordering of allocations in the cycle for a given system is determined by the relative pairwise ratios of all the demands in it. In the example above, $ds^{*}$ is an integer multiple of the other $ds$, with the ratio $2/1$. According to this and the tie breaking policy mentioned above, the cycle consists of $1$ iteration for the user with the $ds^{*}$, followed by $2$ iterations for the user with the lower $ds$.
In a system of $n$ users, the length of the cycle and the number of occurrences for each individual demand are given by
$\lvert l\rvert=\sum_{i=1}^{n}\frac{\operatorname{lcm}DS}{ds_{i}}$
where $\lvert l\rvert$ denotes the length of the cycle, $\operatorname{lcm}DS$ denotes the least common multiple of the set of all dominant shares, $DS$, and $ds_{i}$ is the dominant share of user $i$. The individual occurrences of each demand within the cycle is given by the $i$-$th$ term of the summation for user $i$. In the cases where all other dominant shares are integer fractions of $ds^{*}$, like in the example above, since $lcmS$ is equal to $ds^{*}$, the latter part is equivalent to the ratio of each $ds_{i}$ to $ds^{*}$, and the length of the cycle is equal to the sum of all these ratios, plus $1$, for the single occurrence of $ds^{*}$.
The ratio of lower demands to the $ds^{*}$ is still useful, in cases where integer fraction assumption does not hold. In those cases, further subcycles emerge, characterised by the single occurrence of $ds^{*}$, followed by a differing number of all other demands. The most frequently occurring subcycle is characterised by a single occurrence of the demand with $ds^{*}$, followed by all other demands occurring for times equal to the integer part of the ratio of their dominant share to $ds^{*}$. That is to say:
$\lvert l^{*}\rvert=\sum_{i=1}^{n}\left\lfloor\frac{ds^{*}}{ds_{i}}\right\rfloor$
where $\lvert l^{*}\rvert$ denotes the length of, what we would like to call, the basic subcycle. The other subcycles are characterised by including an additional occurrence of one or more of the lower dominant share demands, the frequency of which is determined by the decimal part of the ratio of $ds^{*}$ to each dominant share, i.e., the $i$-$th$ term of the expression above for demand $i$. Trivially, the first subcycle will always have an extra occurrence, if the ratio has a decimal part.
Whenever a multiple of the decimal part reaches or exceeds an integer number, another allocation for the given demand shows up in the corresponding subcycle. For example, in a system with two demands, whose dominant share ratio is $1.25$, the basic subcycle will consist of each demand taking $1$ allocation, and every $4th$ subcycle will include one more allocation for the lower dominant share, making it $1$ for the highest dominant share and $2$ for the lower. In cases where the decimal summations exceed an integer, such as in the case of a ratio of $1.4$, there will be $2$ different periods of extra occurrence; in this case, after the first subcycle, every $2nd$ and then $3rd$ subcycle iterations.
Systems with more than two users, with all pairwise dominant share ratios having decimal parts, and not necessarily as nice as the ones presented in the examples above, lead to a cornucopia of subcycles, with differing and overlapping extra occurrence periods for each non-highest dominant share demand. Luckily, we will not need to identify all of these for the purposes pursued in the subsequent sections.
3.4 Precomputed Dominant Resource Fairness
Unlike EDRF, PDRF utilises absolute demands, rather than the normalised fractional demands throughout the computation process. For this reason, we start by redefining $d_{ir}$ notation as the absolute demand of user $i$ for resource $r$. PDRF basically computes the resources allocated in one cycle, and then divides the resource vector, $R$, by the resulting sum, the minimum of which gives the number of maximum available iterations $k$, for the cycle:
$k=\min_{r\in R}\left\{\frac{r}{\sum_{i}\left(\frac{ds^{*}}{ds_{i}}\cdot d_{ir}\right)}\right\}=\min_{r\in R}\left\{\frac{r}{ds^{*}\cdot\sum_{i}\frac{d_{ir}}{ds_{i}}}\right\}=\frac{1}{ds^{*}}\cdot\min_{r\in R}\left\{\frac{r}{\sum_{i}\frac{d_{ir}}{ds_{i}}}\right\}$
The individual allocations for each user, in turn, are given by:
$u_{ir}=\left\lfloor k\cdot\frac{ds^{*}}{ds_{i}}\right\rfloor\cdot d_{ir}$
for user $i$. The parentheses enclosing the first two terms denotes the floor function, in order to account for the indivisibility of resources and demands. In the availability of well-divisibility assumption, the algorithm may be relaxed simply by dropping them.
Note that the $1/ds^{*}$ in $k$ and $ds^{*}$ in the second term cancel out nicely to simplify the terms to the following expressions:
$k^{\prime}=\min_{r\in R}\left\{\frac{r}{\sum_{i}\left(\frac{d_{ir}}{ds_{i}}\right)}\right\}$
and
$u_{ir}=\left\lfloor\frac{k^{\prime}}{ds_{i}}\right\rfloor\cdot d_{ir}$
but presumably for numerical reasons, such a simplification leads to less accurate results, and surprisingly, poorer algorithm performance. Thus we do not implement this simplification.
Algorithm 1 Precomputed DRF
1:$R=\left\langle r_{1},r_{2},r_{3},\dots,r_{m}\right\rangle$triangleright resource reserves
2:$C=\left\langle c_{1},c_{2},c_{3},\dots,c_{m}\right\rangle$triangleright consumed amounts of resources
3:$S=\left\langle s_{1},s_{2},s_{3},\dots,s_{n}\right\rangle$triangleright set of dominant shares
4:$U_{i}=\left\langle u_{i1},u_{i2},u_{i3},\dots,u_{im}\right\rangle$triangleright allocation vector for the user i
5:$D_{i}=\left\langle d_{i1},d_{i2},d_{i3},\dots,d_{im}\right\rangle$triangleright demand vector for the user i
6:
7:pick user with the highest dominant share $\text{{pick} user with the highest dominant share }ds^{*}$
8:for $ds_{i}\in DS$ do
9:     $d_{total}\leftarrow d_{total}+\frac{ds^{*}}{ds_{i}}\cdot D_{i}$triangleright Cumulate the total size of demand in a subloop
10:end for
11:
12:$k=\min_{r}\left\{\frac{r}{d_{total}}\right\}$triangleright find maximum available iterations of the subloop
13:
14:for $ds_{i}\in DS$ do
15:     $U_{i}\leftarrow U_{i}+\left\lfloor k\cdot\frac{ds^{*}}{ds_{i}}\right\rfloor\cdot D_{i}$triangleright allocate each user accordingly, for a given subloop
16:     $C\leftarrow C+\left\lfloor k\cdot\frac{ds^{*}}{ds_{i}}\right\rfloor\cdot D_{i}$triangleright update the consumed resources vector
17:end for
18:return
The pseudocode of PDRF is presented in Algorithm 1. A sample implementation can be accessed at this github repository.
4 Results
In this section we present the results of the tests we have carried out by samples drawn from the discrete uniform distribution. $1000$ users are assigned random demands, and resource reserves are likewise assigned randomly. We tested for various numbers of resource types, demand intervals and resource reserve intervals. Here we are presenting a sample which we consider representative. We run each test $1000$ times, and collected the averages, standard deviations, and maximums, presented in Table 1 and Table 2. The scripts that we used to run the tests can be accessed at the same repository, here.
The results show that the deviations of the PDRF allocation from the DRF allocation are limited to a small number of tasks per user, predominantly by $1$, and very rarely exceeding $2$. However, the error is two tailed, meaning, the deviations may either be underallocations or overallocations, the latter of which is a rather serious drawback, since, unlike underallocations, they cannot be compensated for with a final iteration of DRF. Nevertheless, the rate at which these occur are negligibly low.
In the tests carried out with resource intervals drawn from $[50,000-100,000]$ interval and demands are drawn from $[1-10]$ interval, on avarage, $47.7\%$ ($\mu=477.623,n=1000$) of the users are underallocated $1$ task, with a standard deviation of $\sigma=86,373$. The figures for demand intervals $[1-20]$ and $[10-20]$ are similarly, $\mu=474.932,\sigma=38.381$ and $\mu=486.521,\sigma=50.526$, respectively. The overallocations are for demand intervals $[1-10]$ and $[1-20]$ are $\mu=0.694,\sigma=2.137$ and $\mu=0.192,\sigma=0.526$, respectively.
1,000 Users, 10 Resources, Resource Reserves Drawn From [50,000 – 100,000] Interval
Underallocations Overallocations
Demand Intervals 1 Task 2 Tasks Tasks > 2 Maximum Avg. St. Dev. 1 Task 2 Tasks Tasks > 2 Maximum Avg. St. Dev.
1 – 10 477.623 0 0 1 477.623 86.373 0.693 0.001 0 2 0.694 2.137
1 – 20 474.932 0 0 1 474.932 38.381 0.191 0.001 0 2 0.192 0.526
10 – 20 486.521 0 0 1 486.521 50.526 0 0 0 0 0 0
Table 1: Average Case Scenarios
An observation at this point is that the overallocations are biased towards the demands with lower dominant shares, especially the demands containing $1$, with demands of $1$ for all resource types being the most falsely favoured. As seen in Table 1, when the lower end of the interval is moved to $10$, overallocations are totally avoided.
In more general terms, for a given demand, as the ratio of maximum dominant share to its dominant share increases, the number of overallocations increase proportionally. This is caused by large demand and resource reserve intervals which lead to increased variances in the distribution of dominant shares, and mitigated with the increasing number of resource types, since it leads to lower variance among dominant shares. The worst configuration we could cook were with $2$ resource types, as seen in Table 2.
The reason for the algorithm’s implicit favouring of the smaller dominant shares is due to the the rather abstract notion of partial cycles, which are reflected in the decimal points of the number of cycle iterations ($k$). If the exact ordering of the cycle were available, those numbers would indicate the fraction of the cycle that would execute from the beginning, e.g. a decimal point of 0.5 would indicate, first half of the allocations would be done and the cycle would terminate. In the absence of the exact ordering of allocations, the users exploit the partial cycles inversely proportional to the ratio of the $ds^{*}$ to their $ds$.
1000 Users, 2 Resources, Resource Reserves Drawn From [50000 – 100000] Interval
Underallocations Overallocations
Demand Intervals 1 Task 2 Tasks Tasks > 2 Maximum Avg. St. Dev. 1 Task 2 Tasks Tasks > 2 Maximum Avg. St. Dev.
1-100 423.517 0 0 1 423.517 36.429 44.689 5.588 0.48 3 50.407 25.814
1-1000 407.902 0 0 1 407.902 26.373 56.541 14.519 12.989 32 83.959 36.376
Table 2: Worse Case Scenarios
5 Performance of PDRF
In the original DRF article Ghodsi et al. (2011) point out that each allocation operation takes $\mathcal{O}(log(n))$ time, since the selection of lowest allocated dominant share may be done with a binary heap. However, the number of allocation operations is determined by the ratio of the available resource reserves to the demand averages. More rigorously, the expected execution time of the algorithm is given by:
$\mathbb{E}[f(n,R,D)]=\min_{r\in R}\left\{\frac{r}{\mu_{dr}}\right\}\cdot\log(n)$
where, $D$ and $R$ denote the set of demands and resources, respectively, and $\mu_{dr}$ is the demand average for resource $r$. This can lead to performance problems in cases where this ratio is relatively high.
The main contribution of PDRF is that in such cases it can reduce the time of execution dramatically, since it operates in linear time to the number of users and resources, i.e. $\mathcal{O}(m,n)$, and is not effected from this ratio. It takes $m\cdot n$ multiplications to find the fractional demands, $(m-1)\cdot n$ comparisons to find the $ds$ of each demand, $n-1$ comparisons to find $ds^{*}$, $n$ divisions for finding the ratio of $ds^{*}$ to each $ds$, $m\cdot n$ scaling of demands by the obtained ratios, $m\cdot(n-1)$ additions for summing up the scaled demands, $m$ divisions of resources by the demand sums, $m-1$ comparisons to find the minimum of the results of the divisions to decide the number of of cycle iterations, and $2\cdot m\cdot n$ operations for the final assignment of resources, all of which are linear time operations.
Although a final allocation round employing DRF may be needed to satisfy underallocated demands and reach Pareto Efficiency, it iterates over $47\%$ of the number of users, on average, and each of them once, since as shown in Section 4, that is the expected number of users that are short by $1$ task to the DRF allocation. Yet better, since all underallocations are by $1$ task, simply allocating $1$ task, starting from the least allocated user and proceeding in the ascending order until demands are exhausted may suffice, rather than a DRF iteration, which takes $2$ heap operations for user (one taking out and one putting back), and an assignment operation for updating the allocated dominant share value.
As such PDRF offers considerable speed up for large scale applications, such as computing clouds, where the ratio of the resource reserves to the client demands are quite large. Since scheduling is a real time process, in which even small, low level speed ups make an important difference, this point is emphasised even more critically. Cross comparative results from real data may reveal a more accurate measurement, but at this time we cannot provide it, since we do not have access to real data.
6 Discussion
In this section we present three discussions that we are inclined to believe are important. The first is about the case of reiterating PDRF by excluding the saturated cycle’s dominant share and finding further, higher order cycles. The second is a small caveat to the Pareto Efficiency property of DRF, in the demonstration of which we use the cycles approach we developed here. Finally, the third is about the weighting mechanisms and weighting policies for DRF and PDRF.
6.1 Higher Order Cycles
Theoretically, a cycle may need to iterate more than once to reach saturation, since:
$\sum_{i=1}^{n}\left\lfloor k\cdot\frac{ds^{*}}{ds_{i}}\right\rfloor\cdot d_{ir}\leq\sum_{i=1}^{n}\left\lfloor k\cdot\frac{ds^{*}}{ds_{i}}\right\rfloor\cdot d_{ir}$
which leads to residual resources, equal to the difference between the two terms of the inequality above, where the inequality is strict. In fact, in the tests we run, we rarely came across this condition, but reiteration of the cycle did not always improve the allocation accurately. It occasionally lead to overassignment, so we excluded it from the final version of the algorithm. Further scrutiny is needed for functionality of these residual cycles.
The saturation of a cycle is the condition in which $k$ value drops below $1$, so no further iterations of the cycle is possible. In skewed distributions, there may be secondary, tertiary, or even higher order cycles, employment of which can contribute to the accuracy of allocation. These cycles can be detected by iteratively removing $ds^{*}$’s from the demand set, until the $k$ value exceeds $1$ again.
Without loss of generality, we may consider the following single resource scenario where the demands are $2$, $4$, and $10$, for users A, B and C, respectively, and the reserve is $35$ units. User C has the $ds^{*}$, so one cycle iteration consists of $1$ iteration for C, $2$ for B, $4$ for A, and it consumes $2\cdot 4+4\cdot 2+10\cdot 1=26$ units of the reserve. After one iteration of the cycle, $9$ units of resource is left and the cycle cannot take another turn. When we exclude User C, who has the $ds^{*}$, what we are left with is a secondary cycle, with now B having the new $ds^{*}$ and taking $1$ unit, and A taking $2$. The cycle takes one turn consuming 8 resources, and terminates.
It should be noted that this method needs further checks and safeguards, since stacking fair allocations on top of each other does not necessarily lead to the fairness of resulting distribution (i.e. multiple cycle iterations are not guaranteed to be composable) in general, due to the fact that, although allocations are expected to grow at the same pace, the demands are not equally satisfied at the saturation of a cycle. Accumulating residues may cause the allocation drift from the intended final outcome, so they should be investigated in more detail before reaching a conclusion.
In the present study we focused on discrete uniform distribution, and since the issues raised in this section are rather dependent on the initial distribution of demands and resource reserves, they are left out of the scope. Further studies may be carried out to address them in the future.
6.2 Pareto Efficiency
Pareto efficieny is defined as the achieving of an allocation, in which no user’s utility can be increased without decreasing another user’s. Although it is claimed in the original paper that DRF is Pareto Efficient, the algorithm presented thereby fails short to achieve it, if only slightly. Ironically, the answer to this little shortcoming lies in the proof itself.
The proof is based on the claim that "progressive filling for DRF is equivalent to the algorithm presented in Figure $1$" [^1] Since Figure $1$ is a scatter plot, we assume what is meant is Algorithm $1$. (Ghodsi et al., 2011, p. 13). Although it is indicated in the preceding paragraph that the progressive filling algorithm eliminates the saturated demands, no removal condition is given in Algorithm $1$, and neither it is mentioned in the sections and examples describing the algorithm.This might be a minor point eluding the authors’ attention or it might have been assumed that the selection process automatically excludes those demands, but the resulting algorithm as such, cannot achieve Pareto Efficiency in, though rare might they be, certain demand distributions and resource reserves.
At the moment we cannot provide an exact mathematical model for the construction of such configurations, but when the $ds^{*}$ is saturated, and there are more resources available for any other lower demand than its number of occurrences in the cycle, meaning though saturated, $ds^{*}$ will come to be the least allocated with the beginning of the next cycle, the algorithm terminates without allocating available resources to the other users.
An example we rather stumbled upon, which led to this analysis, consists of the same demand vectors in the original article, user A with the demand vector $\left\langle 1,4\right\rangle$ and user B with $\left\langle 3,1\right\rangle$, and the available capacities $\left\langle 59,19\right\rangle$. In such a setting, after 10 iterations, the algorithm halts allocating $\left\langle 2,8\right\rangle$ and $\left\langle 24,8\right\rangle$, respectively, and $\left\langle 33,3\right\rangle$ resources still available, which is sufficient for $3$ more allocations for user B, but since the least allocated dominant share belongs to A, and his demand is saturated, the algorithm cannot proceed. A short simulator script with this example can be found at the same github repository.
As in EDRF, this problem may easily be overcome by excluding the saturated demand from the demand set.
6.3 Weighting Policies
A weighting policy, assigns weights to users, and allocates the resources in proportion to these weights. Fairness, in turn, is defined over the allocations scaled by these weights. In Ghodsi et al. (2011), the question of weighting is addressed briefly, pointing to the fact that it suffices for such a generalisation to assign a weight vector to each user, $W_{i}=\left\langle w_{i1},w_{i2},w_{i3},\dots,w_{im}\right\rangle$, and compute the dominant shares with $ds_{i}=\max_{j}\left\{u_{ij}/w_{ij}\right\}$. In Parkes et al. (2015) study, weights are further defined to be normalised, such that $\sum_{i}w_{ir}=1$, for each resource $r$, and they use another "$\rho$" variable, at the construction of which the weights scale the normalised demands, to be used in their calculations. Again, it is notationwise involved and will not be elaborated on in more detail for the sake of simplicity.
In order to implement a weighting policy for PDRF, it suffices to scale the demands with weights for the calculation of dominant shares, as in DRF. On the other hand, as with EDRF, PDRF needs normalised weights, since the selection of $ds^{*}$, and the proportioning of each $ds$ to it necessitates a standardised measure among them.
At this point, we believe that DRF and its derivatives offer a richer functionality than previously expressed. Even without weights, DRF is an implicit pricing mechanism, implicitly assigning value to resources on the basis of demand and supply dynamics. Inclusion of weights articulates it even stronger. For example, allowing users to purchase resources, or open-endedly invest in resources, and assigning their fraction of the total investment as the user’s weight, DRF can implement an implicit auctioning system. Further scrutinising this issue may be a research topic on its own, and is beyond the scope of the present study, thus we will suffice with pointing to it.
7 Conclusion
The allocation of multiple resources has become an increasingly important problem in the last decade and a half. The seminal study of Ghodsi and his colleagues has created a small literature of its own, since DRF brought many novelties and as much possibilities into this field. The study of Parkes and his colleagues contributes to this literature dearly, by providing a more rigorous theoretical framework to study the properties of DRF in higher resolution. The present study is one situated in this literature, offering yet another analysis method, and an improvement to the original algorithm to reach the same allocation in fewer steps.
One particular contribution of PDRF is that it may be easier to adapt to blockchain context. In fact, it was during our attempt to do so, we came up with the idea, and it lead to the development of a new algorithm for the original context, almost by accident. At that point we turned to a more careful review of the literature, to find EDRF, and the whole endeavour lead to the present study. We are currently working on the final touches of the implementation of the blockchain adaptation, which is a forthcoming study.